OpenStudy (anonymous):

What is the distance between points (2, 8) and (7, 5)? Round to the nearest tenth of a unit. ____UNITS

OpenStudy (igreen):

\(\sf d = \sqrt{(x_2-x_1)^2 + (y_2-y_1)^2}\) (2, 8), (7, 5) x1 y1 x2 y2

OpenStudy (igreen):

Plug the numbers in and simplify.

OpenStudy (anonymous):

@welshfella

OpenStudy (anonymous):

@sammixboo

OpenStudy (igreen):

\(\sf d = \sqrt{(7-2)^2 + (5-8)^2}\) 7 - 2 = ? 5 - 8 = ?

OpenStudy (anonymous):

7-2 is 5 5-8 is 3

OpenStudy (igreen):

-3*

OpenStudy (igreen):

So we have: \(\sf d = \sqrt{(5)^2 + (-3)^2}\) Now simplify 5^2 and -3^2, do you know how?

OpenStudy (anonymous):

ummm nah D:

OpenStudy (igreen):

Okay, here's the thing with exponents. The small number at the top tells us how many times to multiply the number at the bottom. \(\sf 6^2 \rightarrow 6 \times 6 \rightarrow 36\) \(\sf 10^4 \rightarrow 10 \times 10 \times 10 \times 10 \rightarrow 10000\) Get it?

OpenStudy (anonymous):

oh yeah i know what XD exponents

OpenStudy (igreen):

Yep, so can you simplify \(\sf 5^2\) and \(\sf -3^2\)?

OpenStudy (anonymous):

do you multiply it i got -225

OpenStudy (igreen):

No, we add it.

OpenStudy (anonymous):

thats makes way more sense, okay i finally got it thankyou so much dude, seriosuly thankyouuu

OpenStudy (igreen):

No problem, after you add it, find the square root and round it, that will be your answer.

OpenStudy (anonymous):

so is the answer 5.8?